Solve the given problem. Find the following and show your complete solution.


a. Degrees of Freedom


b. Critical Value


c. Confidence Interval


GIVEN:


A sample of six college basketball player had an average weight of 217 pounds with a sample standard deviation of 10 pounds. Find the 95% confidence interval of the true mean weight of all basketball player.


1
Expert's answer
2021-05-04T11:28:09-0400

We have a Sample of 6 basketball players,

Mean, "\mu = 217"

Sample standard deviation, "s = 10"

a.) Degrees of freedom = Sample size - 1 "= 6-1 = 5"

b.) Critical value for 95% confidence interval = 1.96

c.)Confidence interval can be calculated using,

"CI = \mu \pm z\dfrac{s}{\sqrt n}"


"= 217 \pm (0.95) \dfrac{10}{\sqrt 6}"

"CI = 217 \pm 3.89"